PDS_VERSION = PDS3 RECORD_TYPE = STREAM OBJECT = TEXT PUBLICATION_DATE = 1997-03-11 NOTE = "MGS RST Instrument Health Report HEA7070A.TXT" END_OBJECT = TEXT END Miscellaney =========== Our Netscape problems with EPIC over the past week apparently resulted from a renaming of JPL's Flight Operations computers. Our Netscape http proxy host, which was thorn.jpl.nasa.gov, became thorn.fltops.jpl.nasa.gov without notice. Trish Priest reported problems using ACTS that had similar symptoms. Engineering Data ================ From at least 4 PM PST yesterday through 8 AM this morning we were unable to access the MGS query server, so can make no report on instrument health. Access was restored within the past hour or so, and our query for yesterday's engineering data is now in progress. The cause and resolution of the interruption are unknown. Results from the query will be incorporated in our next instrument health report. Open Loop Data ============== Data Collection --------------- No new data Analyses -------- Using ad hoc techniques, we have been able to detect the same 209 second spikes in Doppler data from the GWE test on 1997/051-052 that were reported earlier by John Armstrong. Closed Loop Data ================ Data Collection --------------- We have obtained the following files from OSCAR and EPIC: ATDF: ODF: 97067C068.DAT;1 IONCAL: IONCAL_94_9703_DOPR.CSP;2 TROPCAL: WEATHER: WEATHER_10_LATEST.TXT;19 WEATHER_40_LATEST.TXT;19 WEATHER_60_LATEST.TXT;19 UTPM: EOP: 970307-eop.long 970307-eop.short